OpenStudy (anonymous):

Find the geometric mean between 4√3 and10√3

OpenStudy (tkhunny):

Have you considered multiplying them together and subsequently finding the square root of the product? 1 * 2 ==> sqrt(2) 3*5 ==> sqrt(15)

OpenStudy (anonymous):

I mean I got 6.92 for 4√3 and 17.3 for 10√3 am I doing this right?

OpenStudy (tkhunny):

Well, with \(4\sqrt{3}\)(about 6.92) and \(10\sqrt{3}\) (about 17.32), you should get somehting BETWEEN those values, off a bit toward the lesser. \(4\sqrt{3}\cdot 10\sqrt{3} = 40*3 = 4*10*3\) \(\sqrt{4*10*3} = 2\sqrt{30}\) (about 10.95)